Please help me to solve this equation 1.) x cubes y raised to 4 over xy squared. 2.) 3a raised to 5 b raised to 6 over a squared b cubes. 3.) 6y cubes over 12y raised to seven. 4.) p raised to -3 over p raised to -7. 5.) (a cubes b raised to -3 c raised to zero) raised to -3.
Answer by Cromlix(4381)

1) x^3y^4/xy^2 = x^2y^2
2)3a^5b^6/a^2b^3 = 3a^3b^3
3) 6y^3/12y^7 = 1/2y^4
4) p^-3/p^-7.5 = p^4.5
5 (a^3b^-3c^0)^-3
(a^-9b^9c^0)
(a^-9 b^9)
Hope this helps.
:-)